# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v2

@@ IV-B Resource Allocation Strategies
TABLE V: Summary of dynamic resource allocation methods.
Method Domain Performance Metrics Cloud ECC Model Comp. Res. Alloc.
DySCo [21] Edge-Cloud Infrastructure Deployment Cost, Delay
ADPRL [142] Collaborative Edge Computing Task Completion Time, Energy Consumption
Das et al. [143] Drone Communication Networks Latency, Scalability, Resource Utilization
FRL [81] Cloud-Edge-Terminal IoT Networks Learning Speed, Adaptation
DSCF [144] Distributed Serverless Edge Clouds Inference Latency, Resource Consumption
CADCL-DTM [91] Rolling Bearings Digital Twin Forgetting Performance, RUL Prediction Accuracy
A2C-DRL [145] Edge-Cloud Environments Reward Value, Task Rejection, Load-balancing
Lu et al. [146] IoV Edge Cloud Quality of Service
IntellectReq [147] Device-Cloud Recommendation Communication Efficiency, Recommendation Accuracy
@@
Tian et al. [148] Things-Edge-Cloud Computing Throughput, Delay Utility
FDRL [84] Cloud-Edge Collaborative Manufacturing Makespan, Energy Consumption
DESKG [149] Edge-Cloud Collaborative PSS Service Efficiency, OEE
DRL-MTCN [150] Cloud-Fog-Edge Multi-Tier Networks Revenue, Revenue-Cost Ratio, Acceptance Rate
Fan et al. [123] Edge-Cloud Cooperative Computing Task Processing Delay
@@
Wang et al. [151] IoT-Based Distribution Grid Convergence Rate, Performance Stability
DaCM [100] Edge-Cloud Computing Long-term Profit, Computation Delay
RL [111] Edge-Cloud Networks Execution Time, Network Usage, Migration Delay, Energy Consumption
DODQ [141] Cloud-Edge Computing Task Completion Time
DA-HADRL [152] SAGINs with Cloud-Edge Computing Energy Consumption, Latency
GRLDO [153] Client-Edge-Cloud Environment User Satisfaction
MARL-SF [146] IoV Edge Cloud Quality of Service
AC-DRL [107] 5G Edge-Cloud Continuum QoE, Energy Consumption
DA-DDQN [128] Vehicle-Edge-Cloud Collaboration Latency, Energy Consumption
DRL-LOA [154] Healthcare Cloud-Edge Networks Power Usage, Latency
MADDPG [74] End-Edge-Cloud in 6G Networks Delay, Energy Consumption
MARL-CA [155] Industrial Edge-Cloud Web 3.0 Communication Rate, Stability
OTO [69] Cloud-Edge Collaborative DRL Convergence Speed, Reward
HDMO [138] Edge-Cloud Collaboration Task Completion Time, Cost, Energy Consumption, System Reliability
LsiA3CS [156] Large-Scale IIoT Task Completion Time, Energy Consumption
RIS-DRL [157] RIS-Aided Cloud-Edge Computing Wireless Capacity, Latency, Energy Consumption
@@
Strategic allocation of computational resources across the edge-cloud continuum remains fundamental for maximizing performance in AI-driven collaborative environments. Accordingly, we explore three pivotal strategies: dynamic resource allocation, collaborative resource management, and reinforcement learning-based optimization, each designed to enhance task execution and system efficiency. By effectively distributing resources, these approaches address the challenges of varying workloads and resource constraints in distributed environments. To illustrate their practical impact, Table V provides a comparative analysis of recent dynamic resource allocation methods in ECCC systems. Notably, the table reveals a wide range of application domains—from IoT networks to vehicular edge clouds—alongside key performance metrics like latency, energy consumption, and quality of service, highlighting the adaptability and effectiveness of these strategies in diverse scenarios.
@@ IV-B1 Dynamic Resource Allocation
We highlight key approaches for optimizing ECCC systems through effective task offloading strategies. Edge-cloud systems leverage dynamic resource allocation to address fluctuating workloads and optimize resource utilization, crucial for handling unpredictable real-world applications [55]. Approaches include dynamic NUMA node selection and core consolidation [21], optimizing service placement across the edge-cloud continuum by considering factors like delay and resource constraints. Dynamic task offloading mechanisms further optimize task placement and scheduling based on latency, energy consumption, and task dependencies [142].
In drone communication networks, dynamic resource allocation considers drone mobility, communication loads, and latency constraints for efficient real-time applications [143]. Similarly, for cloud-edge-terminal IoT networks, collaborative policy learning using federated reinforcement learning (FRL) enables dynamic task scheduling, adapting to changing workload demands [81]. Distributed serverless edge clouds utilize dynamic split computing frameworks to minimize inference latency while adhering to resource constraints by dynamically determining splitting points and maintaining container instance warm status [144]. Deep reinforcement learning techniques optimize task offloading decisions in 5G edge-cloud environments, maximizing quality-of-experience (QoE) while meeting latency requirements [107]. Additionally, multi-target-aware dynamic resource scheduling algorithms in multi-tier computing networks consider diverse QoS requirements, including computing, storage, bandwidth, and delay, to enhance resource flexibility and efficiency [150].
@@ IV-B2 Collaborative Resource Management
Efficient ECCC requires coordinating resource allocation across edge nodes and the cloud. Such coordination is essential given the heterogeneous nature of edge devices and dynamic workload demands. For instance, in IoT-based distribution grids, resource optimization across cloud, edge, and end devices spans multiple timescales, with decisions at each level influencing others [151], necessitating sophisticated, interdependent resource management strategies. Dynamic resource allocation is also crucial for adapting to fluctuating demands and optimizing real-time performance [151].
One approach leverages optimal transport (OT) and FL principles. Specifically, an OT-based offline optimization model can initially allocate bandwidth and computation resources, followed by online optimization using federated actor-critic learning (OTFAC) [77]. In this dual-driven approach, edge servers can act as local aggregators within the federated learning process, improving efficiency and reducing delays [160]. Additionally, collaborative resource management benefits from a hybrid approach where edge servers not only collaborate with the cloud but also expand their local resources [100]. This strategy allows edge servers to dynamically purchase cloud resources and expand local capacity for long-term cost optimization [100]. Consequently, this adaptability manages dynamic workloads and varying demands. Edge-cloud and edge-edge cooperation further enhance performance through efficient network-wide resource utilization [123].
@@ IV-B3 Reinforcement Learning for Resource Management
In dynamic edge-cloud systems, RL provides a powerful mechanism for optimizing resource allocation. By interacting with the edge-cloud environment and receiving rewards for performance-enhancing actions, RL agents learn adaptive strategies to improve metrics such as latency [110], energy consumption [111], and task completion time [142]. Specifically, these agents adapt to changing workloads and resource availability, leading to more efficient resource utilization. For example, RL optimizes task offloading decisions in vehicular edge-cloud computing (VECC) to minimize latency while maintaining quality of service [110], and addresses online dependent task offloading (ODTO) in collaborative edge computing by jointly optimizing network flow scheduling and resource allocation [142]. In addition, RL dynamically adjusts model parameters and offloading decisions, improving the adaptability and performance of edge-cloud systems [152]. Researchers have demonstrated RL’s effectiveness in various scenarios, including optimizing microservice placement [111], managing resources for delay-sensitive tasks [110], and handling dependent and parallel tasks [141]. Integrating RL with techniques like graph neural networks (GNNs) [153] and deep Q-networks (DQNs) [141] further enhances the agents’ ability to capture complex relationships and optimize offloading decisions. Multi-agent RL extends these capabilities by enabling decentralized learning and coordinated resource management across multiple servers [145].
@@ IV-C Energy Efficiency Optimization
Power consumption is a critical concern in AI-driven ECCC. We explore strategies for minimizing this consumption without sacrificing performance, focusing on power consumption modeling in Sec. IV-C1, energy-aware scheduling (Sec. IV-C2), and joint optimization techniques in Sec. IV-C3. Sec. IV-C1 provides the foundation for understanding energy dynamics by investigating power consumption modeling within the edge-cloud ecosystem. Building upon these models, Sec. IV-C2 presents energy-aware scheduling algorithms to optimize task allocation and execution for minimal energy use. Subsequently, Sec. IV-C3 explores joint optimization techniques, recognizing the interdependence of energy efficiency and other performance metrics, such as latency, to achieve a balanced and sustainable system operation.
IV-C1 Power Consumption Modeling
Developing energy-efficient edge-cloud systems requires accurate power consumption models that capture power dynamics at both the device and system levels, considering the relationship between workloads and energy consumption. At the device level, models must account for diverse hardware components and their individual power characteristics, adapting to the heterogeneity of edge devices, which range from resource-constrained IoT sensors to powerful edge servers [11]. For example, [161] introduces a comprehensive energy model for the Arm Cortex-M0, considering factors like frequency and wait states that influence energy consumption. Specialized models, such as those in [162] for video transcoding and deep learning on SoC-Clusters, offer insights into workload-specific power dynamics.
System-level modeling encompasses individual device power usage and the energy costs of communication and data transfer between edge, cloud, and edge servers. Dynamic workload fluctuations and task offloading further complicate this process, requiring models to consider the interplay between computation, communication, and storage demands. Research like [163] explores energy-aware task migrations, while [102] addresses resource allocation in heterogeneous environments, optimizing power control and offloading to minimize energy consumption. Incorporating renewable energy and pricing, as in [140], allows for further optimization, while understanding the limitations of cloud-centric approaches in edge environments, as highlighted in [133], is crucial for developing effective models.
IV-C2 Energy-Aware Scheduling
Optimizing the energy efficiency of ECCC hinges on dynamic resource management, including adjusting hardware parameters and employing energy minimization strategies. Dynamic voltage and frequency scaling (DVFS) [164] enables real-time adaptation of processor voltage and frequency according to workload, achieving energy savings by lowering these parameters during periods of low activity. For example, DVFO [164] co-optimizes DVFS and offloading decisions to minimize energy consumption on edge devices during DNN inference. Additionally, sleep-wake mechanisms [165] offer further optimization by selectively powering down idle components. A Co-HDRL algorithm [165] demonstrates this in RIS-aided HetNets by jointly optimizing sleep and power control. Green computing strategies [133] broaden the scope of energy efficiency by considering hardware choices, software optimization, and power management policies. Specifically, approaches like two-phase immersion cooling and energy-centric federated management, explored in [133], contribute to sustainable edge infrastructure.
IV-C3 Joint Optimization Techniques
We examine recent work integrating collaborative intelligence in edge-cloud architectures across diverse application domains. Resource management in edge-cloud systems necessitates a holistic approach, jointly optimizing multiple objectives to balance conflicting goals like minimizing energy consumption and latency. Offloading tasks to the cloud reduces energy consumption on edge devices but introduces latency due to data transfer [102]. Multi-objective optimization (MOO) addresses this by finding Pareto optimal solutions representing the best trade-offs [104]. For example, [19] proposes a multi-objective resource scheduling model considering time, cost, load balance, user satisfaction, and trust, while [166] frames service deployment and task offloading as a multi-task, high-dimensional MOO problem, incorporating network reliability and load balancing. Joint optimization dynamically adjusts power consumption based on real-time workload demands, improving energy efficiency without sacrificing performance [117]. Specifically, [116] proposes an energy-efficient resource allocation scheme for local-edge-cloud collaboration, optimizing latency, resource allocation, and energy consumption using an alopex-based differential evolution algorithm. Similarly, [167] develops a hybrid multi-objective optimization algorithm for resilience in error-prone environments, balancing energy and reliability while meeting quality of service standards. In addition, [140] designs an edge-cloud collaboration system for mixed workloads, aiming to meet delay requirements while reducing carbon emissions and improving profits.
@@ V-A Data Privacy Protection
Robust mechanisms are essential for safeguarding sensitive information throughout its lifecycle in the collaborative edge-cloud paradigm. We explore methods and techniques addressing this critical challenge, ensuring data protection within distributed environments. We begin by examining privacy-preserving learning methods, such as federated learning and secure aggregation, enabling collaborative model training without direct data sharing. Subsequently, we delve into secure multi-party computation (SMPC), which allows joint function computation on private inputs. Additionally, differential privacy techniques, adding noise to protect individual data points while preserving aggregate insights, are considered. Finally, data anonymization approaches, transforming data to prevent individual identification, are explored.
@@ V-A1 Privacy-preserving learning methods
To enhance privacy in FL, techniques like differential privacy (DP) and secure aggregation are employed [168]. DP adds carefully calibrated noise to model updates before aggregation, making it difficult to infer individual data points [169]. Meanwhile, secure aggregation uses cryptographic protocols like multi-party computation (MPC) to aggregate updates without revealing individual contributions to the server [170]. Additionally, zero-concentrated differential privacy (zCDP) offers tighter privacy guarantees than traditional DP in FL [171], and CaPC learning combines MPC, homomorphic encryption, and privately aggregated teacher models to achieve both confidentiality and privacy [172]. Lightweight encryption and aggregation techniques further enhance security and efficiency in federated learning systems [173].
@@ V-A2 Secure Multi-Party Computation
Multiple parties can collaboratively compute a function over their private inputs without revealing them through SMPC [174]. SMPC protocols ensure each party learns only the function’s output, safeguarding data privacy and security, crucial for edge-cloud collaborative AI model training. For example, multiple hospitals could train a joint model on patient data without sharing raw data due to privacy regulations [172]. Additionally, SMPC can decompose complex AI tasks into smaller sub-tasks distributed across edge devices and the cloud, improving training efficiency [175]. Robust management frameworks, such as FlexSMC [176], address the dynamic edge-cloud environment by facilitating node discovery, establishing trust, and ensuring robustness against failures. Coded computation within SMPC, as in coded-MPC (CMPC) [177], further enhances performance through resource optimization, with adaptive gap entangled (AGE) polynomial codes [178] as a promising approach. Ensuring security against malicious attacks during MPC-enabled privacy-preserving training is also addressed by approaches like the one in [179], which provides active security even with a majority of malicious parties. Systems like Helen [180] demonstrate practical applications of maliciously secure cooperative learning for linear models.

### v4

@@ IV-B Resource Allocation Strategies
TABLE V: Summary of resource allocation methods.
Type Method Domain Metrics MC RA ECC
Deep Reinforcement Learning A2C-DRL [13] Edge-Cloud Environments Reward Value, Load-balancing
@@
DA-HADRL [147] SAGINs with Cloud-Edge Energy, Latency
GRLDO [148] Client-Edge-Cloud Environment User Satisfaction
MADDPG [86] End-Edge-Cloud in 6G Networks Delay, Energy Consumption
Federated Learning FRL [93] Cloud-Edge-Terminal IoT Learning Speed, Adaptation
FDRL [95] Cloud-Edge Manufacturing Makespan, Energy Consumption
MARL-SF [149] IoV Edge Cloud Quality of Service
MARL-CA [150] Industrial Edge-Cloud Web3 Communication Rate, Stability
Optimization-Based DySCo [24] Edge-Cloud Infrastructure Deployment Cost, Delay
@@
HDMO [136] Edge-Cloud Collaboration Time, Cost, Energy, Reliability
DaCM [151] Edge-Cloud Computing Long-term Profit, Delay
Model-Aware DSCF [152] Distributed Serverless Edge Inference Latency, Res. Consump.
DRL-LOA [153] Healthcare Cloud-Edge Power Usage, Latency
Wang et al. [154] IoT-Based Distribution Grid Convergence Rate, Stability
Domain-Specific Das et al. [155] Drone Communication Networks Latency, Scalability, Resource Util.
DA-DDQN [139] Vehicle-Edge-Cloud Latency, Energy Consumption
@@
Notes: RA = Resource Allocation.
Strategic allocation of computational resources across the edge-cloud continuum remains fundamental for maximizing performance in AI-driven collaborative environments. Accordingly, we explore three pivotal strategies: dynamic resource allocation, collaborative resource management, and reinforcement learning-based optimization, each designed to enhance task execution and system efficiency. By effectively distributing resources, advanced allocation approaches address the challenges of varying workloads and resource constraints in distributed environments. To illustrate their practical impact, Table V provides a comparative analysis of recent dynamic resource allocation methods in ECCC systems.
Notably, the table reveals a wide range of application domains, from IoT networks to vehicular edge clouds, alongside key performance metrics like latency, energy consumption, and quality of service, highlighting the adaptability and effectiveness of these strategies in diverse scenarios.
Resource allocation strategies in ECCC systems can be systematically organized according to their adaptation mechanisms and coordination approaches: (1) Temporal adaptation strategies (dynamic resource allocation) that respond to time-varying workloads and resource availability through real-time monitoring and reactive adjustment mechanisms; (2) Spatial coordination strategies (collaborative resource management) that optimize resource distribution across geographically distributed edge nodes and cloud infrastructure through inter-node cooperation and hierarchical coordination; (3) Learning-based optimization strategies (reinforcement learning approaches) that continuously improve resource allocation decisions through environmental feedback and experience accumulation. The classification framework emphasizes the complementary nature of these approaches, where temporal adaptation handles immediate demand fluctuations, spatial coordination optimizes cross-infrastructure efficiency, and learning-based methods provide long-term performance improvement through adaptive policy refinement.
@@ IV-B1 Dynamic Resource Allocation
Adaptive resource management addresses fluctuating workloads and optimizes resource utilization through real-time monitoring and adjustment mechanisms, crucial for handling unpredictable application demands [64]. Approaches include dynamic NUMA node selection and core consolidation [24], optimizing service placement across the edge-cloud continuum by considering factors like delay and resource constraints. Dynamic task offloading mechanisms further optimize task placement and scheduling based on latency, energy consumption, and task dependencies [158].
In drone communication networks, dynamic resource allocation considers drone mobility, communication loads, and latency constraints for efficient real-time applications [155]. Similarly, for cloud-edge-terminal IoT networks, collaborative policy learning using federated reinforcement learning (FRL) enables dynamic task scheduling, adapting to changing workload demands [93]. Distributed serverless edge clouds utilize dynamic split computing frameworks to minimize inference latency while adhering to resource constraints by dynamically determining splitting points and maintaining container instance warm status [152].
Deep reinforcement learning techniques optimize task offloading decisions in 5G edge-cloud environments, maximizing quality-of-experience (QoE) while meeting latency requirements [146]. Additionally, multi-target-aware dynamic resource scheduling algorithms in multi-tier computing networks consider diverse QoS requirements, including computing, storage, bandwidth, and delay, to enhance resource flexibility and efficiency [159].
@@ IV-B2 Collaborative Resource Management
Coordinated resource allocation across distributed infrastructure requires sophisticated management strategies that account for heterogeneous device capabilities and dynamic workload demands. For instance, in IoT-based distribution grids, resource optimization across cloud, edge, and end devices spans multiple timescales, with decisions at each level influencing others [154], necessitating sophisticated, interdependent resource management strategies. Dynamic resource allocation is also crucial for adapting to fluctuating demands and optimizing real-time performance [154].
One approach leverages optimal transport (OT) and FL principles. Specifically, an OT-based offline optimization model can initially allocate bandwidth and computation resources, followed by online optimization using federated actor-critic learning (OTFAC) [89]. In such dual-driven approach, edge servers can act as local aggregators within the federated learning process, improving efficiency and reducing delays [160]. Additionally, collaborative resource management benefits from a hybrid approach where edge servers not only collaborate with the cloud but also expand their local resources [151]. Such strategies allow edge servers to dynamically purchase cloud resources and expand local capacity for long-term cost optimization [151]. Consequently, such adaptability manages dynamic workloads and varying demands. Edge-cloud and edge-edge cooperation further enhance performance through efficient network-wide resource utilization [126].
@@ IV-B3 Reinforcement Learning for Resource Management
Adaptive optimization mechanisms leverage reinforcement learning to optimize resource allocation by learning strategies through environmental feedback and performance-enhancing actions [161]. Specifically, these agents adapt to changing workloads and resource availability, leading to more efficient resource utilization. For example, RL optimizes task offloading decisions in vehicular edge-cloud computing (VECC) to minimize latency while maintaining quality of service [161], and addresses online dependent task offloading (ODTO) in collaborative edge computing by jointly optimizing network flow scheduling and resource allocation [158]. In addition, RL dynamically adjusts model parameters and offloading decisions, improving the adaptability and performance of edge-cloud systems [147]. Researchers have demonstrated RL’s effectiveness in various scenarios, including optimizing microservice placement [162], managing resources for delay-sensitive tasks [161], and handling dependent and parallel tasks [143]. Integrating RL with techniques like graph neural networks (GNNs) [148] and deep Q-networks (DQNs) [143] further enhances the agents’ ability to capture complex relationships and optimize offloading decisions. Multi-agent RL extends these capabilities by enabling decentralized learning and coordinated resource management across multiple servers [13].
@@ IV-C Energy Efficiency Optimization
Sustainable power management across edge-cloud environments requires integrated optimization strategies balancing energy consumption with performance through comprehensive modeling and adaptive scheduling mechanisms [163, 11]. Energy-aware scheduling algorithms deliver dynamic adaptation via techniques like DVFS and intelligent sleep-wake management, with DVFO achieving 40% energy savings through co-optimization while Co-HDRL demonstrates 50-70% power reductions in IoT scenarios [164, 165]. Furthermore, joint optimization addresses fundamental trade-offs between efficiency and competing metrics through multi-objective approaches employing Pareto methods to balance conflicting goals [22, 117]. Consequently, advanced strategies demonstrate substantial impact, with green computing achieving 30-45% reductions in data center consumption through renewable integration and intelligent load balancing [142, 122]. Comprehensive analysis is detailed in LABEL:sec:supp-energy of the Supplementary Material.
@@ V Privacy and Security Enhancement
Privacy and security mechanisms in ECCC systems form an interconnected protection ecosystem where techniques often complement and build upon each other rather than operating in isolation. Our classification organizes these mechanisms according to their protection scope and technical approach: (1) Data-centric privacy protection (Sec. V-A) encompassing techniques that protect raw data and sensitive information throughout the distributed learning process, including privacy-preserving learning, secure multi-party computation, differential privacy, and data anonymization; (2) Model-centric security protection (Sec. V-B) addressing threats to AI models themselves, including integrity protection, adversarial attack mitigation, secure updates, and trust management; (3) Communication-centric security (Sec. V-C) securing data transmission and inter-node communication across the edge-cloud continuum.
It is important to note that many privacy and security techniques are inherently non-orthogonal and often complement each other in practical deployments. For example, privacy-preserving learning methods frequently incorporate multiple approaches: differential privacy for statistical privacy guarantees, secure multi-party computation for cryptographic protection, and data anonymization for additional privacy layers. Similarly, federated learning serves dual purposes as both a model adaptation technique (Sec. III-B) and a privacy protection mechanism, demonstrating the multi-faceted nature of techniques in ECCC systems. Understanding these interdependencies is crucial for designing comprehensive security frameworks that leverage synergies between different protection mechanisms while avoiding redundant overhead.
@@ V-A Data Privacy Protection
Comprehensive protection mechanisms safeguard sensitive information throughout its lifecycle in collaborative computing paradigms, addressing critical challenges of data security within distributed environments. We begin by examining privacy-preserving learning methods, such as federated learning and secure aggregation, enabling collaborative model training without direct data sharing. Subsequently, we delve into secure multi-party computation (SMPC), which allows joint function computation on private inputs. Additionally, differential privacy techniques, adding noise to protect individual data points while preserving aggregate insights, are considered. Finally, data anonymization approaches, transforming data to prevent individual identification, are explored.
@@ V-A1 Privacy-preserving learning methods
Privacy enhancement in federated learning leverages techniques including differential privacy (DP) and secure aggregation to protect sensitive information [166]. DP adds carefully calibrated noise to model updates before aggregation, making it difficult to infer individual data points [167]. Meanwhile, secure aggregation uses cryptographic protocols like multi-party computation (MPC) to aggregate updates without revealing individual contributions to the server [168]. Additionally, zero-concentrated differential privacy (zCDP) offers tighter privacy guarantees than traditional DP in FL [169], and CaPC learning combines MPC, homomorphic encryption, and privately aggregated teacher models to achieve both confidentiality and privacy [170]. Lightweight encryption and aggregation techniques further enhance security and efficiency in federated learning systems [171].
@@ V-A2 Secure Multi-Party Computation
Collaborative computation frameworks enable multiple parties to jointly compute functions over private inputs without revealing sensitive data through secure multi-party computation protocols [172]. SMPC protocols ensure each party learns only the function’s output, safeguarding data privacy and security, crucial for edge-cloud collaborative AI model training. For example, multiple hospitals could train a joint model on patient data without sharing raw data due to privacy regulations [170]. Additionally, SMPC can decompose complex AI tasks into smaller sub-tasks distributed across edge devices and the cloud, improving training efficiency [173]. Robust management frameworks, such as FlexSMC [174], address the dynamic edge-cloud environment by facilitating node discovery, establishing trust, and ensuring robustness against failures. Coded computation within SMPC, as in coded-MPC (CMPC) [175], further enhances performance through resource optimization, with adaptive gap entangled (AGE) polynomial codes [176] as a promising approach. Ensuring security against malicious attacks during MPC-enabled privacy-preserving training is also addressed by approaches like the one in [30], which provides active security even with a majority of malicious parties. Systems like Helen [177] demonstrate practical applications of maliciously secure cooperative learning for linear models.